OpenStudy (anonymous):

helpppp: find the critical points: s(t)= (t-1)^4 (t+5)^3

OpenStudy (anonymous):

@mathcalculus , what have you tried?

OpenStudy (anonymous):

yes, i used the product rule: \[\frac{ d }{ dx } (t-1)^{4}* (t+5)^{3}= (t-1)^{4} * 3(t+5)^{2} + (t+5)^{3} * 4(t-1)^{3}\]

OpenStudy (anonymous):

then from there, i'm stuck.

OpenStudy (anonymous):

ok, so far so good. Now, factor out the greatest common factor of both terms. For example, they both have a factor of (t-1)^3, but that's not all they have in common. What else?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Don't they both have a factor of t+5 to a power?

OpenStudy (anonymous):

oh yes, but are from opposite sides..

OpenStudy (anonymous):

doesn't matter the order. AB+CA=A(B+C)

OpenStudy (anonymous):

can you explain to me. i dont really see where we are going with factoring

OpenStudy (anonymous):

The purpose in factoring is to simplify the derivative into a product of three factors. One of the factors will be t-1 to a power, a second factor will be t+5 to a power, and then you will be left with some factor containing a t as well.

OpenStudy (anonymous):

And then once you have your derivative in factored form, you can set each factor equal to zero to find your critical points.

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

i understand what youre saying but not how sure how to do that.

OpenStudy (anonymous):

Ok, lets look at the t-1 term in each expression.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the largest exponent they share?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

now i know that i write (t-1)^3

OpenStudy (anonymous):

but what happens to the 4?

OpenStudy (anonymous):

Exactly, so, we may factor out a \[\left( t-1 \right)^{3}\] from both expressions. Hang on, we will come to that.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Let's take a looke at the first expression in its entirety. If all you are removing is the \[\left( t-1 \right)^{3}\] You would be left with \[4\left( t+5 \right)^{3}\]

OpenStudy (anonymous):

Do you see that?

OpenStudy (anonymous):

no if i was to see the common factor.. i would get (t-1)^3+ (t+5)^2.... then what do i do with the numbers that are left...?

OpenStudy (anonymous):

ok,, there is your mistake. It isn't a sum. It's a product.

OpenStudy (anonymous):

ok so i dont see how you got 4(t+5)^3

OpenStudy (anonymous):

Hang on, I was only talking about factoring out the t-1 from the first expression,, not from the whole thing.

OpenStudy (anonymous):

well you showed me that

OpenStudy (anonymous):

Hang on my pc is giving me fits.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):